A wire has a current density of 7.43 Times 10^6 A/
Answer #1

J = 7.43 * 10^6 A/m^2
A = 1.99 * 10^-6 m^2

We know,
J = I/A
I = J*A
I = 7.43 * 10^6 A/m^2 * 1.99 * 10^-6 m^2
I = 14.78 A

I = Q/t

Charge of 1 electron = 1.6 * 10^-19 C
Charge of 5.87 * 10^20 electrons, = 1.6 * 10^-19 * 5.87 * 10^20 = 93.92 C

t = Q/I
t = 93.92/14.78
t = 6.35 s
Time, t = 6.35 s